# Distributed implementation of tensor-product selected configuration interaction

### v1

@@ Abstract
In recent years, the hybrid “QC+HPC” strategy—where quantum computers screen important determinants, followed by exact diagonalization on classical computers—has shown great potential in the study of strongly correlated systems in quantum chemistry. Last year, an IBM team proposed a novel scheme that utilizes quantum computers to select important bit strings that are then used to construct a spin-adapted determinant space via tensor products. Inspired by this, we have specifically designed a completely new algorithm for this tensor-product selected configuration interaction (SCI). Moreover, for the first time worldwide, we have implemented distributed storage of the CI vector in an SCI program, enabling efficient handling of large-scale computation. Since this study is independent and does not involve determinant selection by quantum computers, we employed our SCI program to conduct full configuration interaction (FCI) computations. Our FCI calculations for N2 (aug-cc-pVDZ) under D2h symmetry and CN (cc-pVTZ) under C2v symmetry, involving $1.47\times 10^{11}$ and $4.86\times 10^{11}$ determinants, respectively, exceed the previous record of $2\times 10^{9}$ determinants computed with the DICE program [J. Chem. Phys. 149, 214110 (2018)] by more than two orders of magnitude. These results set a new benchmark for SCI computations and lay the groundwork for further advancements in SCI methods.
@@ 1 Introduction
The Full Configuration Interaction (FCI) is an exact method for solving the Schrödinger equation in quantum chemistry. 1, 2, 3, 4 However, due to its exponential scaling with system size, FCI remains limited to small molecules with small basis sets. By the use of a distributed CI vector, the largest system calculated to date is propane with the STO-3G basis set, involving up to $1.31\times 10^{13}$ determinants 4.
In practical applications, the coupled cluster singles and doubles with perturbative triples [CCSD(T)] method 5 is known as the gold standard for weakly correlated systems owing to its balanced accuracy and efficiency. While achieving accurate calculations for strongly correlated systems remains the most formidable challenge in quantum chemistry, the selected configuration interaction (SCI) method has long been employed and is recognized as one of the most promising methods for addressing this issue. As early as 1973, Huron et al. introduced the Configuration Interaction using a Perturbative Selection made Iteratively (CIPSI) method 6 in which candidate determinants are selected by rigorously estimating their energy contributions via second-order perturbation theory. Although this selection scheme is considered reliable, its high computational cost precludes its application to larger molecules. Subsequently, methods such as Adaptive Sampling CI (ASCI) 7, 8, and Semistochastic Heat-Bath CI (SHCI) 9, 10, 11 were developed, which can essentially be regarded as approximations to CIPSI.
In recent years, with the rise of quantum computing (QC), several research groups have attempted to rapidly screen important determinants using quantum devices. However, the noise inherent in quantum computers can lead to the loss of some determinants, preventing the generation of wavefunctions that strictly correspond to the eigenstates of the $\hat{S}^{2}$ operator. To address this issue, last year a team from IBM proposed an innovative strategy: using quantum computers to select important bit strings, and then constructing the complete determinants via tensor products. 12 In this paper, we refer to this variant of SCI,which is characterized by this structure, as Tensor-Product SCI (TPSCI).
Since CI calculations are variational, including more determinants generally improves accuracy; however, the associated memory and computational costs for diagonalization increase dramatically. Currently, the widely used DICE program, 9, 10 which distributes the nonzero Hamiltonian matrix elements across nodes, enables calculations with up to $2\times 10^{9}$ determinants. Nevertheless, DICE employs replicated storage for the CI vector, so that the memory available on a single node becomes a bottleneck, thereby limiting the maximum feasible CI vector size despite multi-node execution.
@@
In this work, we have, for the first time worldwide, implemented distributed storage of the CI vector in an SCI program, overcoming single-node memory limitation and enabling the use of the entire supercomputer’s available memory for larger-scale calculations. Moreover, we have designed a completely new, highly efficient algorithm specifically for TPSCI, leveraging its unique tensor-product structure, and further optimized the program for efficient parallel execution on the supercomputer Fugaku. We also validate the use of single precision to reduce computational cost and memory use; this is a strategy pursued for Configuration Interaction as early as the work of Pakiari and Handy, 13 and in line with current trends in correlated methods. 14 The combination of low memory footprint and low precision operations make our work promising for application to leadership class supercomputers based on GPUs.
Ultimately, we computed the FCI energies of N2 (aug-cc-pVDZ) and CN (cc-pVTZ), which, under D2h and C2v symmetry, respectively, involve $1.47\times 10^{11}$ and $4.86\times 10^{11}$ determinants. These results significantly exceed the previously reported maximum SCI scale of $2\times 10^{9}$ determinants computed with the DICE program. 11 These achievements lay a solid foundation for the future development and application of TPSCI. Meanwhile, as previous FCI calculations for N2 15 and CN 16 were performed with the smaller cc-pVDZ basis set, our results serve as valuable benchmark data for further research.
2 CI Vector Distributed Algorithm for TPSCI
In this section, we will describe the parallel algorithm we have designed for TPSCI, including the important optimizations required for large-scale application.
2.1 Distributed CI Vector Storage
In the framework proposed by IBM, only important bit strings are selected, which means that the chosen $\alpha$ and $\beta$ bit strings are identical. The determinant space constructed via tensor products can thus be used to generate closed-shell singlet wavefunctions that strictly correspond to the eigenstates of the $\hat{S}^{2}$ operator.
@@ 2.1 Distributed CI Vector Storage
In our program, the selected $\alpha$ and $\beta$ bit strings are allowed to differ. Suppose their sets are defined as $\{S_{w}^{\alpha}\mid w=1,2,\dots,x\}$ and $\{S_{u}^{\beta}\mid u=1,2,\dots,y\}$, where $S^{\alpha}$ and $S^{\beta}$ are the sets of selected $\alpha$-spin and $\beta$-spin bit strings, respectively, with elements denoted as $S_{w}^{\alpha}$ and $S_{u}^{\beta}$. The determinant space constructed via tensor products is then given by:
$\{|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle\mid w=1,\dots,x;\,u=1,\dots,y\}.$ (1)
@@
$|D_{K}\rangle=|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle,\quad w=\frac{K}{y}+1,\quad u=\text{MOD}(K,y).$ (2)
@@
Since the determinants are indexed such that $\alpha$ bit strings vary first, followed by $\beta$ bit strings, we define a segment as the subset of the CI vector corresponding to all determinants sharing the same $\alpha$ bit string. In this work, we denote the total number of processes (each node runs exactly one process) as $N$. Let $m_{p}$ denote the number of $\alpha$ bit strings assigned to process $p$ (where $p=0,1,\dots,N-1$), with $\sum^{N-1}_{p=0}m_{p}=x$. Then, each process contains $m_{p}$ segments and thus is responsible for a total of $m_{p}\times y$ determinants. As each determinant is fully defined by its index through Eq. (2), our program does not require additional memory to store determinant composition information.
Figure 1: Distribution of the CI vector into segments across $N$ processes. Each process holds a strict subset of $m_{p}$ of the $\alpha$ strings. On a local process, the determinants are divided into segments (represented by vertical arrows), where a segment is composed of all determinants that share the same $\alpha$ string. For this figure, we define a displacement index $d_{p}=\sum_{i=0}^{p-1}m_{i}$ to show the offset of $\alpha$ strings on a given process.
@@ 2.2 Distributed Matrix-Vector Multiplication
In this work, we employ the traditional Davidson diagonalization method 17 for CI calculations, where the core computational step is evaluating $\mathbf{W}=H\cdot\mathbf{U}$, with Hamiltonian matrix $H$, and CI vectors $\mathbf{W}$ and $\mathbf{U}$.
The most advanced SCI implementations to date follow two main strategies: some store the nonzero Hamiltonian elements in a distributed manner across nodes (e.g., DICE), while others compute the Hamiltonian matrix elements on-the-fly (e.g., QUANTUM 18). However, both approaches store the CI vector entirely on a single node without exception. The advantage of these approaches is that each computing node performs a simple matrix-vector multiplication; at the end of each Davidson iteration, the effective Hamiltonian matrix is efficiently constructed using an MPI_ALLREDUCE operation. However, a key limitation is that the CI vector size is strictly constrained by the memory of a single node, posing a major bottleneck in utilizing supercomputing resources for larger determinant spaces.
In our program, the CI vector is distributed across nodes, and each node runs exactly one process. In this setting, each node $p$ only stores its corresponding CI vector segments. The computation of $\mathbf{W_{p}}$ is then given by:
@@
This implies that, in a single Davidson iteration, node $p$ theoretically needs to sequentially fetch data from all other nodes to compute $\mathbf{W_{p}}$. We define each individual fetch-and-compute operation (i.e., computing $\mathbf{W_{p}}\mathrel{+}=H\cdot\mathbf{U_{q}}$) as one step. Since there are $N$ computing nodes, each Davidson iteration consists of $N$ steps. Unlike the implementation in QUANTUM, we precompute and store the diagonal elements of the Hamiltonian matrix in memory, while the off-diagonal elements are computed on-the-fly during the Davidson iterations.
@@
2.3 Efficient Hamiltonian Computation for TPSCI
@@
The Hamiltonian matrix element $\langle D_{w}|H|D_{u}\rangle$ can be expressed as $\langle S_{w}^{\alpha}|\langle S_{w}^{\beta}|H|S_{u}^{\alpha}\rangle|S_{u}^{\beta}\rangle$. According to the Slater-Condon rules, this matrix element can be nonzero only if:
$\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})+\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})\leq 2,$ (4)
where $\text{DIFF}(S_{w},S_{u})$ denotes the number of electron differences between bit strings $S_{w}$ and $S_{u}$. To classify these terms concisely, we introduce the notation $[a,b]$, where $a=\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})$ and $b=\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})$. For example, the term [1,1] corresponds to $\text{DIFF}(S_{w}^{\alpha},S_{u}^{\alpha})=1$ and $\text{DIFF}(S_{w}^{\beta},S_{u}^{\beta})=1$. Throughout this work, we use the shorthand notations [2,0], [1,1], [1,0], [0,2], [0,1], and [0,0] to represent a total of six terms in which the matrix elements may be nonzero. The first five terms correspond to off-diagonal elements of the Hamiltonian matrix, which are always computed on-the-fly. The last term, [0,0], corresponds to diagonal elements, which, although also computable on-the-fly, are precomputed and stored in memory in our program to improve efficiency.
@@
2      for $S_{u}^{\alpha}$leftarrow $S^{\alpha}_{q,1}$ to $S^{\alpha}_{q,m_{q}}$ do
3           case $\leftarrow$ DIFF($S_{w}^{\alpha}$, $S_{u}^{\alpha}$)
@@
5                for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
6                     $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
@@ 7                     Calculate [2,0] term
12                          for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
13                               foreach $S_{u}^{\beta}$in SINGLE_LINK(:,$S_{w}^{\beta}$do
@@ 14                                    Calculate [1,1] term
17                                   $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
@@ 18                                    Calculate [1,0] term
23                                         for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
24                                              foreach $S_{u}^{\beta}$in SINGLE_LINK(:,$S_{w}^{\beta}$do
@@ 25                                                   Calculate [0,1] term
28                                                  $S_{u}^{\beta}\leftarrow S_{w}^{\beta}$
@@ 29                                                   Calculate [0,0] term
36                                                  $S_{u}^{\alpha}$leftarrow $S_{w}^{\alpha}$
37                                                   for $S_{w}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{y}$ do
38                                                        foreach $S_{u}^{\beta}$in the q-th batch of DOUBLE_LINK(:,$S_{w}^{\beta}$do
@@ 39                                                             Calculate [0,2] term
Algorithm 1 Computation algorithm for TPSCI on local process $p$ at step $q$. This algorithm describes how to compute $\langle D_{w}|H|D_{u}\rangle$, where $|D_{w}\rangle$ and $|D_{u}\rangle$ are determinants from processes $p$ and $q$, respectively.
@@
Our novel algorithm for TPSCI is shown in Algorithm 1. Each Davidson iteration consists of $N$ steps. In step $q$, the local process $p$ performs the computation using data fetched from process $q$, which is detailed in Algorithm 1.
@@
To eliminate unnecessary computations, we introduce the SINGLE_LINK and DOUBLE_LINK arrays, which store precomputed single and double excitation connections among $\{S_{w}^{\beta}\mid u=1,2,\dots,y\}$. These arrays ensure that Hamiltonian evaluations only occur when the electron difference between $|D_{w}\rangle$ and $|D_{u}\rangle$ does not exceed 2, directly eliminating invalid cases and improving efficiency. This optimization is enabled due to the unique structure of TPSCI: for each $\alpha$ bit string, the associated $\beta$ bit string set remains unchanged across determinants, allowing excitation connections to be precomputed and reused efficiently.
As the DOUBLE_LINK array can be extremely large, it is distributed nearly evenly across all processes based on its indices. To keep track of this distribution, an auxiliary array is introduced to record the starting and ending indices of DOUBLE_LINK assigned to each process. Consequently, the computation of the [0,2] term is performed batch by batch over the $N$ steps (as shown in Lines 27-32 of Algorithm 1). Since the calculation of the [0,2] term is performed only in the loop over $S_{w}^{\alpha}$, each process $p$ will fetch data from the target process $q$ for the DOUBLE_LINK only once per Davidson step.
Roughly speaking, the computational cost of TPSCI scales as $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, where $N_{\mathrm{SCI}}$ is the number of selected determinants, $N_{\mathrm{occ}}$ the number of occupied orbitals, and $N_{\mathrm{vir}}$ the number of virtual orbitals. This can be further reduced by considering the sparsity of excitation links. The sparsity of single excitation links roughly scales as $\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$, in which $N_{\mathrm{FCI}}$ is the FCI dimension, and thus the overall scaling is more precisely given by $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot N_{\mathrm{SCI}}/N_{\mathrm{FCI}}$.
2.4 Strategies for Efficient Parallel Execution in TPSCI

### v2

# The First Selected Configuration Interaction Implementation with Distributed Storage of the Configuration Interaction Vector

@@ Abstract
[…]
@@ 1 Introduction
Full Configuration Interaction (FCI) is an exact method for solving the Schrödinger equation in quantum chemistry. 1, 2, 3, 4, 5 However, due to its exponential scaling with system size, FCI remains limited to small molecules with small basis sets. Enabled by distributed storage of the configuration interaction (CI) vector 2, as of 2024 FCI implementations have reached a scale of $1.31\times 10^{12}$ determinants. 5
In practical applications, the coupled cluster singles and doubles with perturbative triples [CCSD(T)] method 6 is widely regarded as the gold standard for weakly correlated systems, owing to its balanced accuracy and efficiency. In contrast, accurately describing strongly correlated systems remains one of the most formidable challenges in quantum chemistry. The selected configuration interaction (SCI) method has long been employed and is recognized as one of the most promising methods for addressing this issue. Over the past decades, several SCI methods have been proposed and developed, such as the Configuration Interaction using a Perturbative Selection made Iteratively (CIPSI), 7 Adaptive Sampling CI (ASCI), 8, 9 Semistochastic Heat-Bath CI (SHCI), 10, 11, 12 and iterative CI (iCI). 13, 14, 15 While distributed storage of the CI vector has been successfully applied in FCI calculations, 2, 3, 5 its implementation in SCI poses significant challenges due to the irregular and dynamically evolving nature of the determinant space. To date, all existing SCI implementations have employed replicated CI vectors, which imposes a strict memory limit on the maximum feasible CI vector size, even when multiple nodes are used. As a consequence, the largest SCI calculation reported to date contains just $2\times 10^{9}$ determinants, far smaller than the $1.31\times 10^{12}$ determinants achieved in state-of-the-art FCI computations.
Meanwhile, rapid progress in quantum computing (QC) has led several research groups to explore hybrid approaches that combine quantum and classical computers for studying strongly correlated systems. In 2023, QunaSys first proposed a strategy that uses quantum computers to screen important Slater determinants, followed by exact diagonalization on classical computers—a method referred to as Quantum-Selected Configuration Interaction (QSCI). 16 A downside to this approach is that the noise in quantum computers may cause some determinants to be missed, making the wavefunction deviate from the eigenstates of the $\hat{S}^{2}$ operator. To overcome this issue, in 2024 an IBM team proposed an extension of the QSCI strategy in which quantum computers are used to select important bitstrings, after which the complete determinants are constructed via tensor products of the selected bitstrings. 17 In this paper, we refer to this approach as Tensor-Product Bitstring SCI (TBSCI), as it is based on the tensor product structure of $\alpha$ and $\beta$ bitstrings.
@@
In this work, we have, for the first time worldwide, implemented distributed storage of the CI vector in an SCI program, overcoming single-node memory limitation and enabling the use of the entire supercomputer’s available memory for larger-scale calculations. Moreover, we have designed an entirely new, highly efficient algorithm specifically for TBSCI, leveraging its unique tensor-product structure, and further optimized the program for efficient parallel execution on the supercomputer Fugaku. We also validate the use of single precision to reduce computational cost and memory use; this is a strategy pursued for Configuration Interaction as early as the work of Pakiari and Handy, 18 and in line with current trends in correlated methods. 19 The combination of low memory footprint and low precision operations make our work promising for application to leadership class supercomputers based on GPUs.
Ultimately, we have used our implementation to compute: the potential energy surfaces (PESs) of N2 (aug-cc-pVDZ) and CN (cc-pVTZ), as well as single-point energies of Cr2 (STO-3G) and N2 (cc-pVTZ) near their equilibrium geometries. Among these systems, the largest case is N2 (cc-pVTZ), involving $2.62\times 10^{12}$ determinants under D2h symmetry with the frozen-core approximation. This scale surpasses the previous SCI record of $2\times 10^{9}$ determinants 12 by more than three orders of magnitude, and even doubles the size of the largest FCI calculation reported to date ($1.31\times 10^{12}$ determinants). 5
At present, our work focuses on implementing distributed storage of the CI vector in the Davidson diagonalization step of SCI——a particularly challenging component in the development of large-scale SCI methods. The subsequent second-order perturbative correction on top of SCI has not yet been implemented in such a distributed manner. As a result, our current program cannot perform a complete SCI calculation. Instead, we demonstrate its capabilities by performing FCI calculations, the results of which can serve as valuable benchmark data for future research in quantum chemistry.
2 A Novel and Highly Efficient Algorithm for TBSCI
In this section, we describe a novel and highly efficient algorithm specifically designed for TBSCI, which combines distributed storage of the CI vector with a suite of MPI optimization strategies to enable trillion-scale determinant calculations on modern supercomputers.
2.1 Distributed Storage of CI Vector
@@
In the general framework of TBSCI, the sets of the selected bitstrings can be defined as $\{S_{w}^{\alpha}\mid w=1,2,\dots,l_{\alpha}\}$ and $\{S_{u}^{\beta}\mid u=1,2,\dots,l_{\beta}\}$, where $\{S^{\alpha}\}$ and $\{S^{\beta}\}$ are the sets of selected $\alpha$-spin and $\beta$-spin bitstrings, respectively, with elements denoted as $S_{w}^{\alpha}$ and $S_{u}^{\beta}$. Here, $l_{\alpha}$ denotes the number of selected $\alpha$ and $l_{\beta}$ the number of selected $\beta$ bitstrings. The determinant space constructed via tensor products is then given by:
$\{|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle\mid w=1,\dots,l_{\alpha};\,u=1,\dots,l_{\beta}\}.$ (1)
@@
$|D_{K}\rangle=|S_{w}^{\alpha}\rangle\otimes|S_{u}^{\beta}\rangle,\quad w=\frac{K}{l_{\beta}}+1,\quad u=\text{MOD}(K,l_{\beta}).$ (2)
For convenience, $|D_{K}\rangle$ will be denoted as $|D_{(w,u)}\rangle$ throughout this work.
@@
Since the determinants are indexed such that $\alpha$ bitstrings vary first, followed by $\beta$ bitstrings, we define a segment as the subset of the CI vector corresponding to all determinants sharing the same $\alpha$ bitstring. In this work, we denote the total number of processes (each node runs exactly one process) as $N$. Let $m_{p}$ denote the number of $\alpha$ bitstrings assigned to process $p$ (where $p=0,1,\dots,N-1$), with $\sum^{N-1}_{p=0}m_{p}=l_{\alpha}$. Then, each process contains $m_{p}$ segments and thus is responsible for a total of $m_{p}\times l_{\beta}$ determinants. As each determinant is fully defined by its index through Eq. (2), our program does not require additional memory to store determinant composition information.
Figure 1: Distribution of the CI vector into segments across $N$ processes. Each process holds a strict subset of $m_{p}$ $\alpha$ strings. On a local process, the determinants are divided into segments (represented by vertical arrows), where a segment is composed of all determinants that share the same $\alpha$ string. For this figure, we define a displacement index $d_{p}=\sum_{i=0}^{p-1}m_{i}$ to show the offset of $\alpha$ strings on a given process.
@@ 2.2 Distributed Matrix-Vector Multiplication
In this work, we employ the traditional Davidson diagonalization method 20 for CI calculations, where the core computational step is evaluating $\mathbf{W}=H\cdot\mathbf{U}$, with Hamiltonian matrix $H$, and CI vectors $\mathbf{W}$ and $\mathbf{U}$.
The most advanced SCI implementations to date follow two main strategies: some store the nonzero Hamiltonian elements in a distributed manner across nodes (e.g., DICE 12), while others compute the Hamiltonian matrix elements on-the-fly (e.g., QUANTUM 21). However, both approaches store the CI vector entirely on a single node without exception. The advantage of these approaches is that each computing node performs a simple matrix-vector multiplication; at the end of each Davidson iteration, the effective Hamiltonian matrix can be efficiently constructed using an MPI_ALLREDUCE operation. However, a key limitation is that the CI vector size is strictly constrained by the memory of a single node, posing a major bottleneck in utilizing supercomputing resources for larger determinant spaces.
In our program, the CI vector is distributed across MPI processes, with exactly one process running on each node. In this setting, each node $p$ only stores its corresponding CI vector segments. The computation of $\mathbf{W}_{p}$ is then given by:
@@
This implies that, in a single Davidson iteration, node $p$ theoretically needs to sequentially fetch data from all other nodes to compute $\mathbf{W}_{p}$. We define each individual fetch-and-compute operation (i.e., computing $\mathbf{W}_{p}\mathrel{+}=H\cdot\mathbf{U}_{q}$) as one step. Since there are $N$ computing nodes, each Davidson iteration consists of $N$ steps. Unlike the implementation in QUANTUM, we precompute and store the diagonal elements of the Hamiltonian matrix across processes in a distributed manner consistent with the CI vector, while the off-diagonal elements are computed on-the-fly during the Davidson iterations.
@@
2.3 Efficient Hamiltonian Computation for TBSCI
@@
Suppose the Hamiltonian matrix element we need to compute is $\langle D_{(w,u)}|H|D_{(x,v)}\rangle$, which can be expressed as $\langle S_{w}^{\alpha}|\langle S_{u}^{\beta}|H|S_{x}^{\alpha}\rangle|S_{v}^{\beta}\rangle$. According to the Slater-Condon rules, this matrix element can be nonzero only if:
$\text{DIFF}(S_{w}^{\alpha},S_{x}^{\alpha})+\text{DIFF}(S_{u}^{\beta},S_{v}^{\beta})\leq 2,$ (4)
where $\text{DIFF}(S_{w},S_{x})$ denotes the number of electron differences between bitstrings $S_{w}$ and $S_{x}$. To classify these terms concisely, we introduce the notation $[a,b]$, where $a=\text{DIFF}(S_{w}^{\alpha},S_{x}^{\alpha})$ and $b=\text{DIFF}(S_{u}^{\beta},S_{v}^{\beta})$. For example, the term [1,1] corresponds to $\text{DIFF}(S_{w}^{\alpha},S_{x}^{\alpha})=1$ and $\text{DIFF}(S_{u}^{\beta},S_{v}^{\beta})=1$. Throughout this work, we use the shorthand notations [2,0], [1,1], [1,0], [0,2], [0,1], and [0,0] to represent a total of six terms in which the matrix elements may be nonzero. The first five terms correspond to off-diagonal elements of the Hamiltonian matrix, which are always computed on-the-fly. The last term, [0,0], corresponds to diagonal elements, which, although also computable on-the-fly, are precomputed and stored in memory in our program to improve efficiency.
@@
2      for $S_{x}^{\alpha}$leftarrow $S^{\alpha}_{q,1}$ to $S^{\alpha}_{q,m_{q}}$ do
3           case $\leftarrow$ DIFF($S_{w}^{\alpha}$, $S_{x}^{\alpha}$)
@@
5                for $S_{u}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{l_{\beta}}$ do
6                     $S_{v}^{\beta}\leftarrow S_{u}^{\beta}$
@@ 7                     Calculate [2,0] term
12                          for $S_{u}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{l_{\beta}}$ do
13                               foreach $S_{v}^{\beta}$in SINGLE_LINK(:,$S_{u}^{\beta}$do
@@ 14                                    Calculate [1,1] term
17                                   $S_{v}^{\beta}\leftarrow S_{u}^{\beta}$
@@ 18                                    Calculate [1,0] term
23                                         for $S_{u}^{\beta}$leftarrow $S^{\beta}_{1}$ to $S^{\beta}_{l_{\beta}}$ do
24                                              foreach $S_{v}^{\beta}$in SINGLE_LINK(:,$S_{u}^{\beta}$do
@@ 25                                                   Calculate [0,1] term
28                                                  $S_{v}^{\beta}\leftarrow S_{u}^{\beta}$
@@ 29                                                   Calculate [0,0] term
36                                                  $S_{x}^{\alpha}$leftarrow $S_{w}^{\alpha}$
37                                                   foreach $S_{u}^{\beta}$in the $q$-th batch of {$S^{\beta}$do
38                                                        foreach $S_{v}^{\beta}$in DOUBLE_LINK(:,$S_{u}^{\beta}$do
@@ 39                                                             Calculate [0,2] term
Algorithm 1 Conceptual algorithm for TBSCI on local process $p$ at step $q$. This algorithm describes how to compute $\langle S_{w}^{\alpha}|\langle S_{u}^{\beta}|H|S_{x}^{\alpha}\rangle|S_{v}^{\beta}\rangle$, where $\langle S_{w}^{\alpha}|\langle S_{u}^{\beta}|$ and $|S_{x}^{\alpha}\rangle|S_{v}^{\beta}\rangle$ are determinants from processes $p$ and $q$, respectively. For clarity, we present the algorithm without the symmetry optimization (Sec. 2.4.2) and reassignment strategy (Sec. 2.4.6).
@@
Our novel algorithm for TBSCI is shown in Algorithm 1. Each Davidson iteration consists of $N$ steps. In step $q$, the local process $p$ performs the computation using data fetched from process $q$, which is detailed in Algorithm 1.
@@
To eliminate unnecessary computations, we introduce the SINGLE_LINK and DOUBLE_LINK arrays, which store precomputed single and double excitation connections among $\{S_{u}^{\beta}\mid u=1,2,\dots,l_{\beta}\}$. These arrays ensure that Hamiltonian evaluations only occur when the electron difference between $|D_{(w,u)}\rangle$ and $|D_{(x,v)}\rangle$ does not exceed 2, directly eliminating invalid cases and improving efficiency. This optimization is particularly effective due to the unique structure of TBSCI: for each $\alpha$ bitstring, the associated $\beta$ bitstring set remains unchanged across determinants, allowing excitation connections to be precomputed and reused efficiently. In practice, these connections are stored as one-dimensional arrays to minimize memory usage, accompanied by auxiliary index arrays that record the starting positions and lengths of each column. For ease of explanation, however, we describe them as two-dimensional arrays throughout this paper.
As the DOUBLE_LINK array can be extremely large due to storing all double excitation connections between all $\beta$ bitstrings, it is divided into batches, one per process, each containing a subset of $\{S_{u}^{\beta}\mid u=1,2,\dots,l_{\beta}\}$. For each $S_{u}^{\beta}$, the corresponding column DOUBLE_LINK(:, $S_{u}^{\beta}$) is constructed and stored entirely on a single process, and an auxiliary array records the index range assigned to each process to facilitate data access. Consequently, the computation of the [0,2] term is carried out batch by batch over the $N$ steps (Lines 27–32 of Algorithm 1). Since this term involves only a loop over $S_{w}^{\alpha}$, without referencing $S_{x}^{\alpha}$, each process $p$ needs to fetch DOUBLE_LINK data from the target process $q$ only once per Davidson step.
In contrast, the SINGLE_LINK array has to be replicated on each node; otherwise, MPI communication would be too frequent, as it is accessed over both $S_{w}^{\alpha}$ and $S_{x}^{\alpha}$ (Line 11 of Algorithm 1). It is reasonable to assume that the sparsities of single and double excitations scales as $\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$. Thus, the storage requirement of SINGLE_LINK is approximately $l_{\beta}\cdot N_{\mathrm{occ}}\cdot N_{\mathrm{vir}}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$, where $N_{\mathrm{occ}}$ and $N_{\mathrm{vir}}$ denote the numbers of occupied and virtual orbitals, respectively. This may become a memory bottleneck when $l_{\beta}$ becomes large; in such a case, single excitations may be generated on-the-fly, at the cost of additional computation. However, since the limitation did not arise in the present study and may not occur in future applications—given that $\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$ can be very small—we do not pursue its resolution here.
As shown in Algorithm 1, this sparsity factor ($\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$) appears once in the computations of the $[2,0]$ and $[0,2]$ terms, but twice in that of the $[1,1]$ term. Consequently, the dominant contributions come from the $[2,0]$ and $[0,2]$ terms, each scaling as $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$, while the $[1,1]$ term has a smaller scaling of $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot N_{\mathrm{SCI}}/N_{\mathrm{FCI}}$. In practice, the $[0,2]$ term is the most time-consuming, as the information for double excitations between $S_{w}^{\beta}$ and $S_{u}^{\beta}$ is evaluated within the innermost loop. As mentioned earlier, our TBSCI program can also be used to perform FCI calculations. In this case, the sparsity factor becomes exactly one, and the computational cost simplifies to $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, with the dominant contributions arising from the $[2,0]$, $[1,1]$, and $[0,2]$ terms. In practice, when our implementation is used for FCI the [1, 1] term is the most time-consuming. It should be emphasized that the computational cost of the Handy–Knowles FCI algorithm scales only as $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}\cdot N_{\mathrm{vir}}$; as such, our intention is not to compete with existing FCI methods, but rather to showcase both the algorithmic effectiveness of the newly designed TBSCI method and the high efficiency enabled by our parallelization strategies. A summary of the scaling behavior and the dominant terms is provided in Table 1.
Table 1: Computational scaling and dominant terms of SCI and FCI calculations in the framework of TBSCI.
Method Overall scaling Dominant terms
SCI $N_{\mathrm{SCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}$ $[2,0]$, $[0,2]$
FCI $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$ $[2,0]$, $[1,1]$, $[0,2]$
2.4 Strategies for Efficient Parallel Execution in TBSCI